# Speeding up an unsteady flow simulationM. HANEK, J. PAPEŽ, and J. ŠÍSTEK

### v1

@@ 3.2 Multilevel BDDC preconditioner
In the standard 2-level BDDC method, the coarse degrees of freedom define a global coarse problem with the unknowns ${\bf u}_{C}$ and local subdomain problems with mutually independent degrees of freedom ${\bf u}_{i}$.
@@
(12) ${\bf z}^{\Gamma}=\sum_{i=1}^{N_{S}}\left({\bf R}_{i}^{\Gamma}\right)^{T}{\bf D}_{i}{\bf R}_{Bi}\left({\bf u}_{i}+{\bf\Phi}_{i}{\bf R}_{Ci}{\bf u}_{C}\right).$
Vectors ${\bf u}_{C}$ and ${\bf u}_{i}$ are obtained in each iteration by solving
(13) ${\bf K}_{C}{\bf u}_{C}=\sum_{i=1}^{N_{S}}{\bf R}_{Ci}^{{T}}{\bf\Phi}_{i}^{{T}}{\bf R}_{Bi}^{T}{\bf D}_{i}{\bf R}_{i}^{\Gamma}{\bf r}^{\Gamma},$
(20) $\left[\begin{array}[c]{cc}{\bf K}_{i}&{\bf C}_{i}^{{T}}\\ {\bf C}_{i}&{\bf 0}\end{array}\right]\left[\begin{array}[c]{c}{\bf u}_{i}\\ {\bf\mu}_{i}\end{array}\right]=\left[\begin{array}[c]{c}{\bf R}_{Bi}^{T}{\bf D}_{i}{\bf R}_{i}^{\Gamma}{\bf r}^{\Gamma}\\ {\bf 0}\end{array}\right]\,,\,i=1,\dots,N_{S},$
where ${\bf K}_{C}$ is the stiffness matrix of the global coarse problem, ${\bf K}_{i}$ is the local matrix assembled from elements in the $i$-th subdomain, and ${\bf C}_{i}$ is a constraint matrix enforcing zero values of the local coarse degrees of freedom in the second block row of (20). The restriction matrix ${\bf R}_{Bi}$ selects the local interface unknowns from those at the whole subdomain, the columns of ${\bf\Phi}_{i}$ contain the local coarse basis functions, and ${\bf R}_{Ci}$ is the restriction matrix of the global vector of coarse unknowns to those present at the $i$-th subdomain.
Matrix ${\bf D}_{i}$ applies weights to satisfy the partition of unity. In this work, it corresponds to a diagonal matrix with entries given either by the inverse cardinality of the set of subdomains sharing the interface unknown (denoted as card) or derived from the diagonal entries of the local stiffness matrices (denoted as diag). These are rather standard choices in the DD literature (see, e.g., [26]).
@@ 3.3 Adaptive selection of coarse degrees of freedom
The idea of adaptive BDDC is to enrich the coarse space by additional degrees of freedom. These are chosen so that the condition number of the preconditioned operator $\kappa({\bf M}_{BDDC}^{-1}{\bf A})$ is reduced in a close-to-optimal way. In the present work, we apply the method described in [16, 18, 15]. It is combined with the three-level BDDC preconditioner as a special case of adaptive multilevel BDDC [24]. Although adaptive BDDC was developed for problems requiring a large number of iterations such as those with jumps in material parameters, the studied problem presents another potential use case in which strengthening the preconditioner at the cost of additional computations in its setup may lead to reduction of the overall number of iterations and, ultimately, the computational time.
@@
(21) ${\bf\Pi}\left({\bf I}-{\bf E}_{st}\right)^{{T}}{\bf A}_{st}\left({\bf I}-{\bf E}_{st}\right){\bf\Pi}{\bf w}=\lambda{\bf\Pi}{\bf A}_{st}{\bf\Pi}{\bf w},$
@@
(22) ${\bf A}_{st}=\left[\begin{array}[]{cc}{\bf A}_{s}&\\ &{\bf A}_{t}\end{array}\right],$
@@
(23) ${\bf c}^{st}_{\ell}={\bf w}_{\ell}^{{T}}{\bf\Pi}\left({\bf I}-{\bf E}_{st}\right)^{{T}}{\bf A}_{st}\left({\bf I}-{\bf E}_{st}\right){\bf\Pi}.$
@@
The part of the vector ${\bf c}^{st}_{\ell}$ that corresponds to the unknowns within a face between the subdomains is added as a new row into the matrices ${\bf C}_{s}$ and ${\bf C}_{t}$ from the respective problems (13). Details of the employed approach can be found in [24].
@@
The prescribed threshold $\tau>1$ is an approximation of the target condition number of the preconditioned system $\kappa({\bf M}_{BDDC}^{-1}{\bf A})$. The smaller the threshold, the more coarse degrees of freedom are added. This has the effect of reducing the number of iterations at the cost of increasing the size of the global coarse problem, and hence the cost of each action of the BDDC preconditioner. Consequently, there is a trade-off between pushing the number of iterations down and the cost of each of them in adaptive BDDC. It should be also noted that although the solution of the local generalized eigenvalue problems is parallelized, it still presents a significant overhead in the setup of the BDDC preconditioner that is realized before solving the problem in the first time step.
4 Stopping criteria based on (algebraic) residual
A common (often the default) criterion to terminate an iterative solver is a prescribed reduction of the relative residual. More precisely, a criterion like
(24) $\frac{\|{\bf r}_{k}\|}{\|{\bf r}_{0}\|}<tol$
is considered, where ${\bf r}_{k}$ is the residual in the $k$-th iteration, ${\bf r}_{k}={\bf b}-{\bf A}{\bf x}_{k}$, and ${\bf r}_{0}$ is the residual associated with the initial guess ${\bf x}_{0}$. The key benefit of this criterion is that it involves only (cheaply) computable quantities. On the other hand, one should be aware of its significant limitations:
  1. 1.
    For ill-conditioned systems, a small residual norm does not guarantee a small norm of the error. Recall the well-known relationship
    $\frac{\|{\bf x}-{\bf x}_{k}\|}{\|{\bf x}\|}\leq\kappa({\bf A})\frac{\|{\bf r}_{k}\|}{\|{\bf b}\|},$
[…]
  2. 2.
    The norm $\|{\bf r}_{k}\|$ depends on the discretization basis and has no continuous counterpart, i.e., it is not equal to a norm of some continuous function.
  3. 3.
    It is not clear how to choose the tolerance $tol$ in (24).
A use of an efficient preconditioner, which is our case, resolves the first point raised above and partially also the second one; preconditioned residual is close to the error for very well-conditioned systems. Nevertheless, setting properly the tolerance is a challenging issue that is a subject of our future work. We expect that it should be based on a posteriori error estimates for errors due to the discretization and linearization of (1). In this work, we use a common choice $tol=10^{-6}$.
There is another good reason why (24) may be a bad choice when solving sequences of linear systems. An accurate approximation ${\bf x}_{0}\approx{\bf x}$ (for example by using the approximation computed for the previous system) typically makes the norm of the initial residual ${\bf r}_{0}$ small. The requirement of further improving the norm of the residual ${\bf r}_{k}$ by several orders of magnitude then may lead to useless iterations and an effect of over-solving; after certain level, the error in computing $\psi^{(n+1)}$ in (3) is dominated by the FEM discretization error and reducing the algebraic error further brings no improvement. Consequently, it is important to normalize $\|{\bf r}_{k}\|$ by a quantity that is not related to the initial approximation. In this study, we take the norm of the right-hand side vector ${\bf b}$ for this purpose, and we terminate the iterations when
(25) $\frac{\|{\bf r}_{k}\|}{\|{\bf b}\|}<10^{-6}.$
It seems essential to combine the choice of the non-zero initial solution ${\bf x}_{0}$ with this stopping criterion in order to achieve the reduction of the number of iterations expected from an accurate initial guess. Especially for problems that converge (with increasing time) to a particular solution and hence the subsequent algebraic systems become progressively closer to each other.
Finally, note that (25) is used as a stopping criterion for PCG for example in MATLAB R2019b, SciPy (all versions), or PETSc 3.21. However, (24) was the default option in PETSc until the 3.18 release, and it might still be used in some other software.
5 Krylov subspace recycling
@@
Deflation brings two challenges: how to construct $\mathcal{U}$, and how to implement the iterative method on $\mathcal{V}$. While the latter is a technical issue that has been resolved for many iterative (Krylov) methods and is independent of the choice of $\mathcal{U}$ (and $\mathcal{V}$), a proper choice of the deflation subspace is problem-dependent and should be carefully addressed for each application.
@@ 5.1 Deflated preconditioned conjugate gradient method
For a square matrix ${\bf A}\in\mathbb{R}^{n\times n}$ and a vector ${\bf v}\in\mathbb{R}^{n}$, define the $j$-th Krylov subspace
$\mathcal{K}_{j}({\bf A},{\bf v})=\mbox{span}\{{\bf v},{\bf A}{\bf v},\ldots,{\bf A}^{j-1}{\bf v}\}.$
Given a symmetric positive definite matrix ${\bf A}$, a right-hand side ${\bf b}$, and an initial approximation ${\bf x}_{0}$, define the initial residual ${\bf r}_{0}={\bf b}-{\bf A}{\bf x}_{0}$. In the $j$-th iteration, CG generates an approximation ${\bf x}_{j}$ characterized by
${\bf x}_{j}\in{\bf x}_{0}+\mathcal{K}_{j}({\bf A},{\bf r}_{0}),\qquad{\bf r}_{j}={\bf b}-{\bf A}{\bf x}_{j}\perp\mathcal{K}_{j}({\bf A},{\bf r}_{0})$
The implementation of CG can be found, e.g., in the seminal paper [12].
Let us store basis vectors of the subspace $\mathcal{U}$ as columns of matrix ${\bf W}$. Define
(26) ${\bf Q}={\bf I}-{\bf W}({\bf W}^{T}{\bf A}{\bf W})^{-1}{\bf W}^{T}{\bf A},$
and, for ${\bf v}$ such that ${\bf W}^{T}{\bf v}=0$,
$\mathcal{K}_{\mathcal{U},j}({\bf A},{\bf v})=\mathcal{U}+\mathcal{K}_{j}({\bf Q}{\bf A}{\bf Q},{\bf v}).$
Here ${\bf Q}$ is a projector on $\mathcal{V}$, which is an orthogonal complement of $\mathcal{U}$ with respect to the inner product induced by matrix ${\bf A}$. Therefore, $\mathcal{K}_{j}({\bf Q}{\bf A}{\bf Q},{\bf v})\cap\mathcal{U}=\{0\}$. Deflated Conjugate Gradient method is a modification of CG that generates approximations such that
@@
assuming that ${\bf x}_{0}$ is such that ${\bf W}^{T}{\bf r}_{0}=0$. A corresponding algorithm is given in [22, Algorithm 3.5]; see also [8, Sect. 3]. For the ease of presentation, we will call ${\bf W}$ the deflation basis.
To speed up the convergence of CG (or any iterative method in general), a suitable preconditioning formally transforming the problem ${\bf A}{\bf x}={\bf b}$ into ${\bf M}^{-1}{\bf A}{\bf x}={\bf M}^{-1}{\bf b}$ is often considered. Preconditioning can be also used in combination with the deflation. The corresponding algorithm is given, e.g., in [22, Algorithm 3.6], and we provide it below as Algorithm 1 for completeness.
@@
Mathematically, it holds that ${\bf W}^{T}{\bf r}_{j}={\bf W}^{T}{\bf p}_{j}=0$. In finite-precision computations, however, this may not hold due to the loss of orthogonality. In such cases, the reorthogonalization of residuals at line 12 of Algorithm 1 may be necessary; see also [22, Eq. (7.1)].
@@ 5.2 Subspace recycling
When solving a sequence of systems ${\bf A}{\bf x}^{(i)}={\bf b}^{(i)}$, the idea might be to reuse, or recycle, the Krylov subspace built for the previous system(s) to construct the deflation space for the current system. We denote the vectors computed in (deflated) PCG applied to ${\bf A}{\bf x}^{(i)}={\bf b}^{(i)}$ by the superscript, for example ${\bf p}^{(i)}_{j}$ denotes the $j$-th search vector for the $i$-th system. The maximal dimension of the deflation space is a parameter to be chosen by the user, we denote it by $R$.
@@
The deflation basis is set as the first $R$ search vectors and then remains unchanged for the rest of the computation, ${\bf W}^{(i)}={\bf W}=[{\bf p}^{(1)}_{1},{\bf p}^{(1)}_{2},\ldots,{\bf p}^{(1)}_{R}]$. If $R$ is larger than the number $\ell$ of PCG iterations for solving the first system, the basis ${\bf W}$ also contains the search vectors from solving the second system (and possibly more systems until it has $R$ columns), ${\bf W}=[{\bf p}^{(1)}_{1},\ldots,{\bf p}^{(1)}_{\ell},{\bf p}^{(2)}_{1},\ldots]$.
@@
The basis is first constructed as in the previous case, and then it uses a ‘sliding window’ to contain the last $R$ search vectors from the previous systems. This choice is motivated by the fact that the search vectors from the preceding linear systems may be more relevant to the subsequent system to be solved. In more detail, after solving the $(i-1)$-th system and saving $\ell$ search vectors as columns of ${\bf P}^{(i-1)}$, the deflation basis is updated as
@@
The search vectors ${\bf p}_{j}$ are in PCG, as well as in deflated PCG ([22, Prop. 3.3]), ${\bf A}$-orthogonal, ${\bf p}_{j}^{T}{\bf A}{\bf p}_{l}=0$ for $j\neq l$. Therefore, if the basis ${\bf W}$ contains the search vectors only, the matrix ${\bf W}^{T}{\bf A}{\bf W}$ is diagonal and the projection ${\bf Q}$ in (26) only consists of matrix multiplications. In finite-precision computations, the ${\bf A}$-orthogonality between the search vectors is often lost during the course of iterations. However, in our application with the BDDC preconditioner, only a decent number of PCG iteration was needed to reach the tolerance, and the loss of orthogonality was modest in our experiments.
@@
(27) ${\bf V}^{T}{\bf A}{\bf V}{\bf y}=\theta{\bf V}^{T}{\bf M}{\bf V}{\bf y}$
and setting ${\bf w}={\bf V}{\bf y}$ as an approximation to an eigenvector of ${\bf M}^{-1}{\bf A}$. This standard approximation, however, requires to apply the preconditioner ${\bf M}$ to ${\bf V}$, which may not be possible when only the operation ${\bf M}^{-1}{\bf v}$ is available, as in our case of the BDDC preconditioner. As in [22, Sect. 5.1], we therefore consider the harmonic Ritz approximation
(28) ${\bf V}^{T}{\bf A}{\bf M}^{-1}{\bf A}{\bf V}{\bf y}=\theta{\bf V}^{T}{\bf A}{\bf V}{\bf y}$
that involves the operation with the inverse of the preconditioner. Moreover, we employ a strategy from [22, Sections 5.1 and 5.2] for constructing matrices ${\bf V}^{T}{\bf A}{\bf M}^{-1}{\bf A}{\bf V}$ and ${\bf V}^{T}{\bf A}{\bf V}$ from (28) for ${\bf V}=[{\bf W}^{(i-1)},{\bf P}^{(i-1)}]$ at a low cost. Notice that while problems (27) and (28) are different, they both provide approximations to the eigenvalues and eigenvectors of the preconditioned system matrix ${\bf M}^{-1}{\bf A}$.
After solving one or several systems, the maximal size $R$ of the deflation basis ${\bf W}^{(i-1)}$ is reached, and only a subset of $R$ harmonic Ritz approximations (eigenpairs satisfying (28) and setting ${\bf w}={\bf V}{\bf y}$) are taken to form ${\bf W}^{(i)}$. We consider two variants:
@@
(29) $\frac{\|\bm{\theta}^{(i)}-\bm{\theta}^{(i-1)}\|_{2}}{\|\bm{\theta}^{(i)}\|_{2}}\leq 10^{-5},$

### v3

@@ 3.2 Multilevel BDDC preconditioner
In the standard 2-level BDDC method, the coarse degrees of freedom define a global coarse problem with the unknowns ${\bf u}_{C}$ and local subdomain problems with mutually independent degrees of freedom ${\bf u}_{i}$. How to obtain ${\bf u}_{C}$ and ${\bf u}_{i}$ will be discussed later.
@@
${\bf z}^{\Gamma}=\sum_{i=1}^{N_{S}}\left({\bf R}_{i}^{\Gamma}\right)^{T}{\bf D}_{i}{\bf R}_{Bi}\left({\bf u}_{i}+{\bf\Phi}_{i}{\bf R}_{Ci}{\bf u}_{C}\right),$ (7)
where the restriction matrix ${\bf R}_{i}^{\Gamma}$ selects the local interface unknowns from the global interface unknowns, ${\bf R}_{Bi}$ selects the local interface unknowns from those at the whole subdomain, the columns of ${\bf\Phi}_{i}$ contain the local coarse basis functions, and ${\bf R}_{Ci}$ is the restriction matrix of the global vector of coarse unknowns to those present at the $i$-th subdomain. Matrix ${\bf D}_{i}$ applies weights to satisfy the partition of unity. In this work, it corresponds to a diagonal matrix with entries given either by the inverse cardinality of the set of subdomains sharing the interface unknown (denoted as card) or derived from the diagonal entries of the local stiffness matrices (denoted as diag). These are rather standard choices in the DD literature; see, e.g., [31].
Vectors ${\bf u}_{C}$ and ${\bf u}_{i}$ in (7) are obtained in each iteration by solving
${\bf K}_{C}{\bf u}_{C}=\sum_{i=1}^{N_{S}}{\bf R}_{Ci}^{{T}}{\bf\Phi}_{i}^{{T}}{\bf R}_{Bi}^{T}{\bf D}_{i}{\bf R}_{i}^{\Gamma}{\bf r}^{\Gamma},$ (8)
$\left[\begin{array}[c]{cc}{\bf K}_{i}&{\bf C}_{i}^{{T}}\\ {\bf C}_{i}&{\bf 0}\end{array}\right]\left[\begin{array}[c]{c}{\bf u}_{i}\\ {\bf\mu}_{i}\end{array}\right]=\left[\begin{array}[c]{c}{\bf R}_{Bi}^{T}{\bf D}_{i}{\bf R}_{i}^{\Gamma}{\bf r}^{\Gamma}\\ {\bf 0}\end{array}\right]\,,\,i=1,\dots,N_{S},$ (15)
where ${\bf K}_{C}$ is the stiffness matrix of the global coarse problem, ${\bf K}_{i}$ is the local matrix assembled from elements in the $i$-th subdomain, and ${\bf C}_{i}$ is a constraint matrix enforcing zero values of the local coarse degrees of freedom in the second block row of (15).
@@
As shown in [33], the condition number of the Schur complement preconditioned by BDDC $\kappa({\bf M}_{BDDC}^{-1}{\bf A})=\lambda_{\max}/\lambda_{\min}\leq C\left(1+\log^{2}(H/h)\right)$, where $H$ is the characteristic subdomain size and $h$ is the characteristic element size. In other words, the condition number is independent of the number of subdomains and grows only mildly with their size. The analysis reveals that the smallest eigenvalue $\lambda_{\min}\geq 1$, and the bound for $\kappa({\bf M}_{BDDC}^{-1}{\bf A})$ is actually the upper bound on the largest eigenvalue $\lambda_{\max}$ [34].
@@ 3.3 Adaptive selection of coarse degrees of freedom
The idea of adaptive BDDC is to enrich the coarse space with additional degrees of freedom. These are chosen to reduce the condition number of the preconditioned operator $\kappa({\bf M}_{BDDC}^{-1}{\bf A})$ in a close-to-optimal way. In the present work, we apply the adaptive multilevel BDDC method described in [29]. Although adaptive BDDC was originally developed for problems requiring a large number of iterations, such as those with jumps in material parameters, the problem studied here represents another potential use case. In this case, strengthening the preconditioner, at the cost of additional setup computations, may reduce the overall number of iterations and, ultimately, the total computational time.
@@
${\bf\Pi}\left({\bf I}-{\bf E}_{st}\right)^{{T}}{\bf A}_{st}\left({\bf I}-{\bf E}_{st}\right){\bf\Pi}{\bf w}=\lambda{\bf\Pi}{\bf A}_{st}{\bf\Pi}{\bf w},$ (16)
@@
${\bf A}_{st}=\left[\begin{array}[]{cc}{\bf A}_{s}&\\ &{\bf A}_{t}\end{array}\right],$ (17)
@@
${\bf c}^{st}_{\ell}={\bf w}_{\ell}^{{T}}{\bf\Pi}\left({\bf I}-{\bf E}_{st}\right)^{{T}}{\bf A}_{st}\left({\bf I}-{\bf E}_{st}\right){\bf\Pi}.$ (18)
@@
The part of the vector ${\bf c}^{st}_{\ell}$ that corresponds to the unknowns within a face between the subdomains is added as a new row into the matrices ${\bf C}_{s}$ and ${\bf C}_{t}$ from the respective problems (8). In this work, arithmetic averages on edges and faces of the subdomains are used as the initial coarse degrees of freedom, and the eigenvectors corresponding to the largest eigenvalues of the generalized eigenvalue problems are used to enrich this set. Details of the employed approach can be found in [29].
@@
The prescribed threshold $\tau>1$ is an approximation of the target condition number of the preconditioned system $\kappa({\bf M}_{BDDC}^{-1}{\bf A})$. With the minimal eigenvalue $\lambda_{\min}=1$ given, the aim is to reduce the maximal eigenvalue $\lambda_{\max}$ according to $\tau$. The smaller the threshold, the more coarse degrees of freedom are added. This has the effect of reducing the number of iterations at the cost of increasing the size of the global coarse problem, and hence the cost of each action of the BDDC preconditioner. Consequently, there is a trade-off between pushing the number of iterations down and the cost of each of them in adaptive BDDC. It should be also noted that although the solution of the local generalized eigenvalue problems is parallelized, it still presents a significant overhead in the setup of the BDDC preconditioner that is realized before solving the problem in the first time step.
For implementation reasons, the number of eigenvectors computed on each face is limited to 10, and the maximal number of iterations of the employed Locally Optimal Block Preconditioned Conjugate Gradient (LOBPCG) method [35] is limited to 15. Due to these limits, the threshold $\tau$ is only a rather weak indicator of the final condition number of the preconditioned system, typically underestimating it significantly.
4 Deflation and Krylov subspace recycling
@@
Deflation brings two challenges: i) how to construct $\mathcal{U}$, and ii) how to implement the iterative method on $\mathcal{V}$. While the latter is a technical issue that has been resolved for many iterative (Krylov) methods and is independent of the choice of $\mathcal{U}$ (and $\mathcal{V}$), a proper choice of the deflation subspace is problem-dependent and should be carefully addressed for each application.
@@ 4.1 Deflated preconditioned conjugate gradient method
Given a symmetric positive definite matrix ${\bf A}\in\mathbb{R}^{n\times n}$, a right-hand side ${\bf b}\in\mathbb{R}^{n}$, and an initial approximation ${\bf x}_{0}$, define the initial residual ${\bf r}_{0}={\bf b}-{\bf A}{\bf x}_{0}$. In the $j$-th iteration, CG ([36]) generates an approximation ${\bf x}_{j}$ characterized by
${\bf x}_{j}\in{\bf x}_{0}+\mathcal{K}_{j}({\bf A},{\bf r}_{0}),\qquad{\bf r}_{j}={\bf b}-{\bf A}{\bf x}_{j}\perp\mathcal{K}_{j}({\bf A},{\bf r}_{0}),$
where $\mathcal{K}_{j}({\bf A},{\bf v})=\mbox{span}\{{\bf v},{\bf A}{\bf v},\ldots,{\bf A}^{j-1}{\bf v}\}$ is the $j$-th Krylov subspace.
Let columns of a matrix $\bf W$} give the basis of the subspace $\mathcal{U}$. Then
${\bf Q}={\bf I}-{\bf W}({\bf W}^{T}{\bf A}{\bf W})^{-1}{\bf W}^{T}{\bf A},$ (19)
is a projector on $\mathcal{V}$, which is an orthogonal complement of $\mathcal{U}$ with respect to the inner product induced by matrix $\bf A$}. For $\bf v$} such that ${\bf W}^{T}{\bf v}=0$ define
$\mathcal{K}_{\mathcal{U},j}({\bf A},{\bf v})\equiv\mathcal{U}+\mathcal{K}_{j}({\bf Q}{\bf A}{\bf Q},{\bf v}).$}
From the properties of $\bf Q$}, $\mathcal{K}_{j}({\bf Q}{\bf A}{\bf Q},{\bf v})\cap\mathcal{U}=\{0\}$.
Deflated Conjugate Gradient method is a modification of CG that generates approximations such that
@@
assuming that ${\bf x}_{0}$ is such that ${\bf W}^{T}{\bf r}_{0}=0$. A corresponding algorithm is given in [17, Algorithm 3.5]; see also [19, Sect. 3]. For the ease of presentation, we will call $\bf W$} the deflation basis.
To speed up the convergence, a suitable preconditioning formally transforming the problem ${\bf A}{\bf x}={\bf b}$ into ${\bf M}^{-1}{\bf A}{\bf x}={\bf M}^{-1}{\bf b}$ is often considered. Preconditioning can be also used in combination with the deflation. The corresponding algorithm is given, e.g., in [17, Algorithm 3.6], and we provide it below as Algorithm 1 for completeness.
@@
Mathematically, it holds that ${\bf W}^{T}{\bf r}_{j}={\bf W}^{T}{\bf p}_{j}=0$. In finite-precision computations, however, this may not hold due to the loss of orthogonality. In such cases, the reorthogonalization of residuals at line 12 of Algorithm 1 may be necessary; see also [17, Eq. (7.1)].
The deflation can be elegantly employed also in the case when matrix $\bf A$} is only symmetric positive semidefinite with a known nullspace. This is the case when problem (3) becomes a pure Neumann problem for $\partial\Omega_{O}=\emptyset$. In this case, the nullspace of $\bf A$} is spanned by constant vectors, and a basis vector of ones can be added as one column of $\bf W$}. Note that the nullspace of Schur complement $\bf A$} is a restriction of the nullspace of $\bf K$} in (5) from all unknowns to the interface $\Gamma$.
@@ 4.2 Subspace recycling
When solving a sequence of systems ${\bf A}{\bf x}^{(i)}={\bf b}^{(i)}$, the idea might be to reuse, or recycle, the Krylov subspace built for the previous system(s) to construct the deflation space for the current system, as in [18]. We denote the vectors computed in (deflated) PCG applied to ${\bf A}{\bf x}^{(i)}={\bf b}^{(i)}$ by the superscript, for example ${\bf p}^{(i)}_{j}$ denotes the $j$-th search vector for the $i$-th system. The maximal dimension of the deflation space is a parameter to be chosen by the user, we denote it by $R$.
@@
The deflation basis is set as the first $R$ search vectors and then remains unchanged for the rest of the computation, ${\bf W}^{(i)}={\bf W}=[{\bf p}^{(1)}_{1},{\bf p}^{(1)}_{2},\ldots,{\bf p}^{(1)}_{R}$]. If $R$ is larger than the number $\ell$ of PCG iterations for solving the first system, the basis $\bf W$} also contains the search vectors from solving the second system (and possibly more systems until it has $R$ columns), ${\bf W}=[{\bf p}^{(1)}_{1},\ldots,{\bf p}^{(1)}_{\ell},{\bf p}^{(2)}_{1}$$\ldots$].
@@
The basis is first constructed as in the previous case, and then it uses a ‘sliding window’ to contain the last $R$ search vectors from the previous systems. This choice is motivated by the fact that the search vectors from the preceding linear systems may be more relevant to the subsequent system to be solved. In more detail, after solving the (i1)(i-1)-th system and saving $\ell$ search vectors as columns of ${\bf P}^{(i-1)}$, the deflation basis is updated as
@@
The search vectors ${\bf p}_{j}$ are in PCG, as well as in deflated PCG ([17, Prop. 3.3]), $\bf A$}-orthogonal, ${\bf p}_{j}^{T}{\bf A}{\bf p}_{l}=0$ for $j\neq l$. Therefore, if the basis $\bf W$} contains the search vectors only, the matrix ${\bf W}^{T}{\bf A}{\bf W}$ is diagonal and the projection $\bf Q$} in (19) only consists of matrix multiplications. In finite-precision computations, the $\bf A$}-orthogonality between the search vectors is often lost during the course of iterations. However, in our application with the BDDC preconditioner, only a decent number of PCG iteration was needed to reach the tolerance, and the loss of orthogonality was modest in our experiments.
@@
${\bf V}^{T}{\bf A}{\bf V}{\bf y}=\theta{\bf V}^{T}{\bf M}{\bf V}{\bf y}$ (20)
and setting ${\bf w}={\bf V}{\bf y}$ as an approximation to an eigenvector of ${\bf M}^{-1}{\bf A}$. This standard approximation, however, requires to apply the preconditioner $\bf M$} to $\bf V$}, which may not be possible when only the operation ${\bf M}^{-1}{\bf v}$ is available, as in our case of the BDDC preconditioner. As in [17, Sect. 5.1], we therefore consider the harmonic Ritz approximation
${\bf V}^{T}{\bf A}{\bf M}^{-1}{\bf A}{\bf V}{\bf y}=\theta{\bf V}^{T}{\bf A}{\bf V}{\bf y}$ (21)
that involves the operation with the inverse of the preconditioner. Moreover, we employ a strategy from [17, Sects. 5.1 and 5.2] for constructing matrices ${\bf V}^{T}{\bf A}{\bf M}^{-1}{\bf A}{\bf V}$ and ${\bf V}^{T}{\bf A}{\bf V}$ from (21) for ${\bf V}=[{\bf W}^{(i-1)},{\bf P}^{(i-1)}]$ at a low cost. Notice that while problems (20) and (21) are different, they both provide approximations to the eigenvalues and eigenvectors of the preconditioned system matrix ${\bf M}^{-1}{\bf A}$.
After solving one or several systems, the maximal size $R$ of the deflation basis ${\bf W}^{(i-1)}$ is reached, and only a subset of $R$ harmonic Ritz approximations (eigenpairs satisfying (21) with ${\bf w}={\bf V}{\bf y}$) are taken to form ${\bf W}^{(i)}$. We consider two variants:
@@
$\frac{\|\bm{\theta}^{(i)}-\bm{\theta}^{(i-1)}\|_{2}}{\|\bm{\theta}^{(i)}\|_{2}}\leq 10^{-5},$ (22)
@@
An interesting approach for selecting the deflation basis with the aid of proper orthogonal decomposition (POD) was presented in [38] in the context of evolving system matrix. The relation between this POD-based deflation and harmonic Ritz vectors, which are studied in the present paper, remains to be investigated.
5 Stopping criteria based on (algebraic) residual
A common (often the default) criterion to terminate an iterative solver is a prescribed reduction of the relative residual as
$\frac{\|{\bf r}_{k}\|}{\|{\bf r}_{0}\|}<tol,$ (23)
where ${\bf r}_{k}$ is the residual in the $k$-th iteration, ${\bf r}_{k}={\bf b}-{\bf A}{\bf x}_{k}$, and ${\bf r}_{0}={\bf b}-{\bf A}{\bf x}_{0}$ is the residual associated with the initial guess ${\bf x}_{0}$. Clearly, this criterion involves only (cheaply) computable quantities. On the other hand, one should be aware of its significant limitations; (23) does not guarantee a small norm of the error, the norm $\|{\bf r}_{k}\|$ depends on the discretization basis, and it is not clear how to choose the tolerance toltol.
There is another good reason why (23) may be a bad choice when solving sequences of linear systems. An accurate approximation ${\bf x}_{0}\approx{\bf x}$ typically makes the norm of the initial residual ${\bf r}_{0}$ small. Such initial approximation might be given as the (final) approximation computed for the previous system or using more elaborated strategies, e.g., based on reduced order models and proper orthogonal decomposition as in [39]. The requirement of further improving the norm of the residual ${\bf r}_{k}$ by several orders of magnitude then may lead to useless iterations and an effect of over-solving; after certain level, the error in computing $\psi^{(n+1)}$ in (3) is dominated by the FEM discretization error and reducing the algebraic error further brings no improvement. Consequently, it is important to normalize $\|{\bf r}_{k}\|$ by a quantity that is not related to the initial approximation. In this study, we take the norm of the right-hand side vector $\bf b$} for this purpose, and we terminate the iterations when
$\frac{\|{\bf r}_{k}\|}{\|{\bf b}\|}<10^{-6}.$ (24)
Finally, note that (24) is used as a stopping criterion for PCG for example in MATLAB R2019b, SciPy (all versions), or PETSc 3.21. However, (23) was the default option in PETSc until version 3.18, and it might still be used in some other software packages.